Use the values in the table below to answer the following questions.

Q

TC

ATC

MC

1

14.52

14.52

5

A

20

6

B

C

51.5

22

4958

225.4

23

E

D

675


2. Cell C represents average total cost= 6 units. what should be the value of cell c?
Answer #1

Answer - The value of C cell will be 25.25.

Value of cell A = 20 *5 = 100

Value of cell B = 100 + 51.5(mc)

= 151.5

Value of cell C = cell B / 6

= 151.5 / 6

= 25 .35